### v5

# On a discrete version of the position-momentum commutation relation

Abstract
The usual position-momentum commutation relation plays a fundamental role in the mathematical description of continuous-variable quantum systems. In the case of a qudit described by a Hilbert space of a high enough dimension, there exists a discrete version of the position-momentum commutation relation satisfied with approximation of a large part of the pure quantum states. Our purpose is to explore in more details the set of these states. We show that it contains a family of discrete-variable Gaussian states depending on a continuous parameter and certain discrete coherent states. It also contains various discrete-variable versions of the Hermite-Gauss states, defined either as eigenstates of certain discrete versions of the harmonic oscillator Hamiltonian or generated by using a discrete version of the creation or annihilation operator. The presented results suggest a possible way to experimentally obtain some of the investigated states.
preprint: AIP/123-QED
I Introduction
The quantum systems with finite-dimensional Hilbert space Schwinger60 ; Vourdas04 ; Cotfas11 ; Cotfas23 ; Cotfas24 play an important role in the investigation of quantum mechanics foundations, in Quantum Information, and in other applications. The mathematical description of these discrete-variable systems is obtained by following, as much as possible, the analogy with the continuous-variable quantum systems. In certain cases, the discrete-variable versions have properties very similar to those concerning the continuous systems, but this does not happen in all the cases.
In quantum mechanics, a pair $A$ and $B$ of Hermitian operators on a Hilbert space is canonical if they satify the relation $A,B]\!=\!\mathrm{i}\hbar$, where [A,B]=ABBA[A,B]\!=\!AB-\!BA, $h\!=\!2\pi\hbar$ is the Plank constant, and $\mathrm{i}\hbar\equiv\mathrm{i}\hbar I$ is the identity operator multiplied by $\mathrm{i}\hbar$. In the finite-dimensional case, such a pair does not exist because the trace of the two sides of $A,B]\!=\!\mathrm{i}\hbar$ is different. In Farrales25 , the authors show that certain canonical pairs can be obtained by relaxing the definition, namely by imposing the relation $A,B]\!=\!\mathrm{i}\hbar$ to be satisfied only on a proper subspace of the Hilbert space. Some references concerning possible applications of this new definition are presented in Farrales25 .
Our approach is different. We show that a discrete version of the position-momentum commutation relation $\bar{q},\bar{p}]\!=\!\mathrm{i}\hbar$ is satisfied with an approximation in a proper subspace if the dimension of the Hilbert space is high enough. Among the elements of such a subspace there are some of the most important states used in the mathematical description of qudits Mehta87 ; Barker ; Nikiforov ; Vilenkin .
The presented results may be used as a starting point of new mathematical formalisms involving only certain particular states of a qudit, by following, for example, the analogy with the theory of continuous-variable Gaussian states used in Quantum Information Ferraro05 ; Wang07 ; Adesso14 .
II Quantum systems with finite-dimensional Hilbert space
A spinless particle with mass, in one dimension, is usually described by using the Hilbert space
$L^{2}(\mathbb{R})\!=\!\overline{\left\{\Psi\!:\!\mathbb{R}\!\rightarrow\!\mathbb{C}\ \left|\ \int_{-\infty}^{\infty}|\Psi(q)|^{2}\,dq\!<\!\infty\right.\right\}}$ (1)
(closure of the space of square integrable functions) with
$\left\langle\Phi,\Psi\right\rangle\!=\!\int_{-\infty}^{\infty}\overline{\Phi(q)}\,\Psi(q)\,dq.$ (2)
In order to have a simple analogy, we shall consider only the case of a discrete-variable quantum system described by an odd-dimensional (d=2s+1d\!=\!2s\!+\!1) Hilbert space $\mathcal{H}$ regarded as a space of functions, namely
$\mathcal{H}\!=\!\{\psi\!:\!\{-\!s,\!-s\!+\!1,...,s\!-\!1,s\}\!\rightarrow\!\mathbb{C}\,\}$ (3)
with the inner product
$\langle\varphi|\psi\rangle\!=\!\sum\limits_{n=-s}^{s}\overline{\varphi(n)}\,\psi(n)$ (4)
and the norm
$\psi||\!=\!\sqrt{\langle\psi|\psi\rangle}\!=\!\sqrt{\sum\limits_{n=-s}^{s}|\psi(n)|^{2}}.$ (5)
The set $\{-\!s,\!-s\!+\!1,...,s\!-\!1,s\}$ contains one and only one representative of any class of remainders modulo $d$. Therefore, we can identify it with $\mathbb{Z}_{d}$. The discrete $\delta$ functions $\delta_{-s}$, $\delta_{-s+1}$, … , $\delta_{s}$, where $\delta_{m}\!:\!\{-\!s,\!-s\!+\!1,...,s\}\!\rightarrow\!\mathbb{C},$
[…]
form an orthonormal basis of $\mathcal{H}$. By using Dirac notation $|m\rangle$ instead of $|\delta_{m}\rangle$, we have
$\langle n|m\rangle\!=\!\delta_{nm},\quad\mbox{and}\quad\sum_{m=-s}^{s}|m\rangle\langle m|\!=\!\mathbb{I},$ (7)
where $\mathbb{I}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H},\ \mathbb{I}\psi\!=\!\psi$, is the identity operator.
II.1 Discrete Fourier transform
The usual Fourier transform $\Psi\!\mapsto F\Psi,$
$\begin{array}[]{l}(F\Psi)(p)\!=\!\frac{1}{\sqrt{2\pi\hbar}}\int\limits_{-\infty}^{\infty}\mathrm{e}^{-\frac{\mathrm{i}}{\hbar}pq}\,\Psi(q)\,dq\!=\!\frac{1}{\sqrt{h}}\int\limits_{-\infty}^{\infty}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{h}pq}\,\Psi(q)\,dq,\end{array}$ (8)
is a unitary transform. Its inverse is the adjoint transform $\Psi\!\mapsto F^{\dagger}\Psi,$
$\begin{array}[]{l}(F^{\dagger}\Psi)(p)\!=\!\frac{1}{\sqrt{h}}\int\limits_{-\infty}^{\infty}\mathrm{e}^{\frac{2\pi\mathrm{i}}{h}pq}\,\Psi(q)\,dq,\end{array}$ (9)
The corresponding discrete-variable version $\mathbf{F}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H}\!:\!\psi\mapsto\mathbf{F}\psi,$
$(\mathbf{F}\psi)(k)\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\!\sum\limits_{n=-s}^{s}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}kn}\,\psi(n),$ (10)
is also a unitary transform Vourdas04 ; Mehta87 . Its inverse is the adjoint $\mathbf{F}^{\dagger}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H}\!:\!\psi\mapsto\mathbf{F}^{\dagger}\psi,$
$(\mathbf{F}^{\dagger}\psi)(k)\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\!\sum\limits_{n=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}kn}\,\psi(n).$ (11)
Particularly, $\langle n|\mathbf{F}|k\rangle\!=\!(\mathbf{F}\delta_{k})(n)\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}kn}$, $\langle n|\mathbf{F}^{\dagger}|k\rangle\!=\!(\mathbf{F}^{\dagger}\delta_{k})(n)\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}kn}$, and
$\begin{array}[]{l}\mathbf{F}\!=\!\mathbb{I}\mathbf{F}\mathbb{I}\!=\!\sum_{n,k=-s}^{s}|n\rangle\langle n|\mathbf{F}|k\rangle\langle k|\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\!\sum\limits_{n,k=-s}^{s}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}kn}\,|n\rangle\langle k|,\\[2.84526pt] \mathbf{F}^{\dagger}\!=\!\mathbb{I}\mathbf{F}^{\dagger}\mathbb{I}\!=\!\sum_{n,k=-s}^{s}|n\rangle\langle n|\mathbf{F}^{\dagger}|k\rangle\langle k|\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\!\sum\limits_{n,k=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}kn}\,|n\rangle\langle k|.\end{array}$ (12)
II.2 Discrete version of the position and momentum operators
The position operator $\hat{q}\!:\!D_{q}\!\subset\!L^{2}(\mathbb{R})\!\rightarrow\!L^{2}(\mathbb{R})\!:\!\Psi\!\mapsto\hat{q}\,\Psi,$
$\begin{array}[]{l}(\hat{q}\,\Psi)(q)\!=\!q\,\Psi(q),\end{array}$ (13)
whose domain of definition is
[…]
admits Schwinger60 ; Vourdas04 the discrete-variable version $\hat{\mathbf{q}}:\mathcal{H}\!\rightarrow\!\mathcal{H}\!:\!\psi\mapsto\hat{\mathbf{q}}\psi,$
$(\hat{\mathbf{q}}\psi)(n)\!=\!n\,\psi(n).$ (15)
The basis $\{|n\rangle\}_{n=-s}^{s}$ is an eigenbasis for $\hat{\mathbf{q}}$, and $\hat{\mathbf{q}}$ admits the spectral decomposition
$\hat{\mathbf{q}}\!=\!\sum\limits_{n=-s}^{s}n|n\rangle\langle n|.$ (16)
The momentum operator $\hat{p}\!:\!D_{p}\!\subset\!L^{2}(\mathbb{R})\!\rightarrow\!L^{2}(\mathbb{R})\!:\!\Psi\!\mapsto\hat{p}\,\Psi,$
$\begin{array}[]{l}\hat{p}\,\Psi\!=\!-\mathrm{i}\hbar\frac{\mathrm{d}}{\mathrm{d}q}\Psi,\end{array}$ (17)
whose domain of definition is
[…]
satisfies the relation
$\hat{p}\!=\!F^{\dagger}\hat{q}F.$ (19)
The discrete-variable version Schwinger60 ; Vourdas04 of momentum is defined as $\hat{\mathbf{p}}\!$$\mathcal{H}\!\rightarrow\!\mathcal{H}$,
$\hat{\mathbf{p}}\!=\!\mathbf{F}^{\dagger}\hat{\mathbf{q}}\mathbf{F}.$ (20)
Explicitly,
$\begin{array}[]{l}(\hat{\mathbf{p}}\psi)(n)\!=\!\frac{1}{d}\sum\limits_{k=-s}^{s}\sum\limits_{m=-s}^{s}k\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}k(n-m)}\,\psi(m),\end{array}$ (21)
and $\hat{\mathbf{p}}$ admits the spectral decomposition
$\hat{\mathbf{p}}\!=\!\sum\limits_{n=-s}^{s}n\,\mathbf{F}^{\dagger}|n\rangle\langle n|\mathbf{F},$ (22)
that is $\{\mathbf{F}^{\dagger}|n\rangle\}_{n=-s}^{s}$ is an eigenbasis for $\hat{\mathbf{p}}$.
II.3 A $d$-dimensional version of Pauli operators
The two-dimensional Pauli operator $Z:\mathbb{C}^{2}\!\longrightarrow\!\mathbb{C}^{2},$
$\begin{array}[]{l}Z|0\rangle\!=\!|0\rangle,\\ Z|1\rangle\!=\!-|1\rangle,\end{array},$ (23)
admits Vourdas04 the general version $\mathbf{Z}\!:\!\mathcal{H}\!\longrightarrow\!\mathcal{H},$
$\mathbf{Z}|n\rangle\!=\!\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}n}|n\rangle$ (24)
satisfying the relations
$\begin{array}[]{l}\mathbf{Z}\!=\!\mathbf{Z}\mathbb{I}\!=\!\mathbf{Z}\sum\limits_{n=-s}^{s}|n\rangle\langle n|\!=\!\sum\limits_{n=-s}^{s}\mathbf{Z}|n\rangle\langle n|\!=\!\sum\limits_{n=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}n}|n\rangle\langle n|,\\[5.69054pt] (\mathbf{Z}\psi)(n)\!=\!\langle n|\mathbf{Z}|\psi\rangle\!=\!\sum\limits_{m=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}m}\langle n|m\rangle\langle m|\psi\rangle\!=\!\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}n}\psi(n),\\[5.69054pt] \mathrm{e}^{\frac{2\pi\mathrm{i}}{d}\hat{\mathbf{q}}}\!=\!\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}\hat{\mathbf{q}}}\mathbb{I}\!=\!\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}\hat{\mathbf{q}}}\sum\limits_{n=-s}^{s}|n\rangle\langle n|\!=\!\sum\limits_{n=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}\hat{\mathbf{q}}}|n\rangle\langle n|\!=\!\sum\limits_{n=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}n}|n\rangle\langle n|\!=\!\mathbf{Z}.\end{array}$ (25)
The two-dimensional Pauli operator $X:\mathbb{C}^{2}\!\longrightarrow\!\mathbb{C}^{2},$
$\begin{array}[]{l}X|0\rangle\!=\!|1\rangle,\\ X|1\rangle\!=\!|0\rangle,\end{array}$ (26)
admits Vourdas04 the general version $\mathbf{X}\!:\!\mathcal{H}\!\longrightarrow\!\mathcal{H},$
$\mathbf{X}|n\rangle\!=\!|n\!+\!1\rangle$ (27)
where the addition is modulo $d$, and we choose $n\!+\!1\!\in\!\{-\!s,\!-s\!+\!1,...,s\!-\!1,s\}$. It has the properties
$\begin{array}[]{l}\mathbf{X}\!=\!\mathbf{X}\mathbb{I}\!=\!\mathbf{X}\sum\limits_{n=-s}^{s}|n\rangle\langle n|\!=\!\sum\limits_{n=-s}^{s}\mathbf{X}|n\rangle\langle n|\!=\!\sum\limits_{n=-s}^{s}|n\!+\!1\rangle\langle n|,\\[5.69054pt] (\mathbf{X}\psi)(n)\!=\!\langle n|\mathbf{X}|\psi\rangle\!=\!\sum\limits_{m=-s}^{s}\langle n|m\!+\!1\rangle\langle m|\psi\rangle\!=\!\psi(n\!-\!1),\\[5.69054pt] \mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}\hat{\mathbf{p}}}\!=\!\mathbf{F}^{\dagger}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}\hat{\mathbf{q}}}\mathbf{F}\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\mathbf{F}^{\dagger}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}\hat{\mathbf{q}}}\sum\limits_{n,k=-s}^{s}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}kn}\,|k\rangle\langle n|\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\mathbf{F}^{\dagger}\sum\limits_{n,k=-s}^{s}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}kn}\,\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}k}|k\rangle\langle n|\\[2.84526pt] \qquad\ \ \ =\!\mbox{\small$\frac{1}{\sqrt{d}}$}\sum\limits_{n,k=-s}^{s}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}k(n+1)}\,\mathbf{F}^{\dagger}|k\rangle\langle n|\!=\!\mbox{\small$\frac{1}{d}$}\sum\limits_{n,k,m=-s}^{s}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}k(n+1)}\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}km}\,|m\rangle\langle n|\\[2.84526pt] \qquad\ \ \ =\!\sum\limits_{n,m=-s}^{s}\mbox{\small$\frac{1}{d}$}\sum\limits_{k=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}k(m-n-1)}\,|m\rangle\langle n|\!=\!\sum\limits_{n,m=-s}^{s}\delta_{m,n+1}\,|m\rangle\langle n|\!=\!\sum\limits_{n=-s}^{s}|n\!+\!1\rangle\langle n|\!=\!\mathbf{X}.\end{array}$ (28)
The operators $\mathbf{Z},\ \mathbf{X}:\!\mathcal{H}\!\longrightarrow\!\mathcal{H}$ are unitary transformations:
$\begin{array}[]{l}\langle\mathbf{Z}\varphi,\mathbf{Z}\psi\rangle\!=\!\sum\limits_{n=-s}^{s}\overline{\mathbf{Z}\varphi(n)}\ \mathbf{Z}\psi(n)\!=\!\sum\limits_{n=-s}^{s}\overline{\varphi(n)}\,\psi(n)\!=\!\langle\varphi,\psi\rangle,\\[2.84526pt] \langle\mathbf{X}\varphi,\mathbf{X}\psi\rangle\!=\!\sum\limits_{n=-s}^{s}\overline{\mathbf{X}\varphi(n)}\ \mathbf{X}\psi(n)\!=\!\sum\limits_{n=-s}^{s}\overline{\varphi(n\!-\!1)}\,\psi(n\!-\!1)\!=\!\langle\varphi,\psi\rangle.\end{array}$ (29)
II.4 A discrete version of the displacement operators
From
$\begin{array}[]{l}\mathbf{Z}\mathbf{X}\!=\!\sum\limits_{n,m=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}n}|n\rangle\langle n|m\!+\!1\rangle\langle m|\!=\!\sum\limits_{m=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}(m+1)}|m\!+\!1\rangle\langle m|,\\[2.84526pt] \mathbf{X}\mathbf{Z}\!=\!\sum\limits_{n,m=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}n}|m\!+\!1\rangle\langle m|n\rangle\langle n|\!=\!\sum\limits_{m=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}m}|m\!+\!1\rangle\langle m|\end{array}$ (30)
it follows that
$\mathbf{Z}\mathbf{X}\!=\!\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}}\,\mathbf{X}\mathbf{Z},$ (31)
and the relations
[…]
In the continuous-variable case, the displacement operators are defined as
$\mathbf{D}(q,p)\!=\!\mathrm{e}^{-\frac{\mathrm{i}}{2\hbar}qp}\,\mathrm{e}^{\frac{\mathrm{i}}{\hbar}p\hat{q}}\,\mathrm{e}^{-\frac{\mathrm{i}}{\hbar}q\hat{p}}\!=\!\mathrm{e}^{-\frac{\pi\mathrm{i}}{h}qp}\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{h}p\hat{q}}\,\mathrm{e}^{-\frac{2\pi\mathrm{i}}{h}q\hat{p}}$. (33)
By following the analogy with the continuous case, the discrete displacement operators are defined as $D(n,k)\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H},$
$\mathbf{D}(n,k)\!=\!\mathrm{e}^{-\frac{\pi\mathrm{i}}{d}nk}\,\mathbf{Z}^{k}\,\mathbf{X}^{n}\!=\!\mathrm{e}^{-\frac{\pi\mathrm{i}}{d}nk}\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}k\hat{\mathbf{q}}}\,\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}n\hat{\mathbf{p}}}$. (34)
They satisfy the relations Vourdas04
$\begin{array}[]{l}\mathbf{D}(n,k)\psi(m)=\mathrm{e}^{-\frac{\pi\mathrm{i}}{d}nk}\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}km}\,\psi(m\!-\!n),\\[5.69054pt] \mathbf{D}(n,k)\,\mathbf{D}(m,\ell)\!=\!\mathrm{e}^{\frac{\pi\mathrm{i}}{d}(km-n\ell)}\ \mathbf{D}(n\!+\!m,k\!+\!\ell)\\[5.69054pt] \mathbf{D}(n,k)\,\mathbf{D}(m,\ell)\!=\!\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}(km-n\ell)}\ \mathbf{D}(m,\ell)\ \mathbf{D}(n,k)\\[5.69054pt] \mathbf{D}^{\dagger}(n,k)\!=\!\mathbf{D}(-n,-k)\!=\!\mathbf{D}^{-1}(n,k).\end{array}$ (35)
III Discrete version of the position-momentum commutation relation
Table 1: Eigenvalues of $\mathbf{C}\!=\![\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}$ in two cases.
Case $d=11$   Case $d=31$
[…]
In the continuous case, the usual position and momentum satisfy the remarkable relation $\hat{q},\hat{p}]={\rm i}\hbar$, that is
$\hat{q},\hat{p}]={\rm i}\mbox{\small$\frac{h}{2\pi}$}.$ (36)
This means that
$\hat{q}(\hat{p}\Psi)-\hat{p}(\hat{q}\Psi)\!=\!{\rm i}\mbox{\small$\frac{h}{2\pi}$}\Psi$ (37)
for any $\Psi$ from the subspace
$\left\{\,\Psi\!\in\!D_{q}\cap D_{p}\ \ |\ \ \hat{q}\Psi\!\in\!D_{p}\ \ \mbox{and}\ \ \hat{p}\Psi\!\in\!D_{q}\,\right\}.$ (38)
If we compare the continuous and discrete version of the Fourier transform, then we arrive at the conclusion that, in the discrete case, the role of Planck constant is played by the dimension $d$ of the Hilbert space. Consequently, the discrete version of the relation (36) is
$\hat{\mathbf{q}},\hat{\mathbf{p}}]={\rm i}\mbox{\small$\frac{d}{2\pi}$}$ (39)
One can see that this relation is not satisfied in the whole $\mathcal{H}$, but by computing the eigenvalues of the operator $\mathbf{C}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H},$
$\begin{array}[]{l}\mathbf{C}\!=\![\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\end{array}$ (40)
with the matrix elements
$\begin{array}[]{l}\langle n|\mathbf{C}|m\rangle\!=\!\frac{1}{d}\sum\limits_{k=-s}^{s}k(n\!-\!m)\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}k(n-m)}-{\rm i}\mbox{\small$\frac{d}{2\pi}$}\delta_{nm}\end{array}$ (41)
for different values of $d$, one arrives at the conclusion that the eigenvalues $\lambda$ with $|\lambda|<0.001$ represent $36\%$ in the case d=11d\!=\!11, represent $64\%$ in the case d=31d\!=\!31, represent $77\%$ in the case d=61d\!=\!61 and $84\%$ in the case d=101d\!=\!101 (see Table I). So, for $d$ large enough, a significant part of the eigenvalues of $\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}$ are almost null.
For any $\varepsilon\!>\!0$, there exists $d_{\varepsilon}$ such that the set
[…]
is non-empty for $d\!=\!\mathrm{dim}(\mathcal{H})\!>\!d_{\varepsilon}$. If $\varepsilon\!>\!0$ is small enough, then we can consider that
$\begin{array}[]{l}[\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\frac{d}{2\pi}\psi\quad\mbox{for}\ \ \psi\!\in\!\mathcal{S}_{\varepsilon}.\end{array}$ (43)
If the states $\psi_{1}$, $\psi_{2}$, … , $\psi_{k}$ from $\mathcal{S}_{\varepsilon}$ form an orthonormal system and $\psi\!=\!\sum\limits_{m=1}^{k}\alpha_{m}\psi_{m}$ is such that $\psi||\!=\!1$, then
[…]
that is $\psi\!\in\!\mathcal{S}_{\varepsilon\sqrt{k}}$.
III.1 A spectral decomposition of the Fourier transform
By using the relation
$(\mathbf{F}\hat{\mathbf{q}}\mathbf{F}^{\dagger})\psi(n)\!=\!\frac{1}{d}\sum\limits_{m,k=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}m(k-n)}\,m\,\psi(k)\!=\!\frac{1}{d}\sum\limits_{m,k=-s}^{s}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}m(k-n)}\,(-m)\,\psi(k)\!=\!-(\mathbf{F}^{\dagger}\hat{\mathbf{q}}\mathbf{F})\psi(n)\!=\!-\hat{\mathbf{p}}\psi(n)$
we get $\mathbf{F}\hat{\mathbf{q}}\mathbf{F}^{\dagger}\!=\!-\hat{\mathbf{p}}$ and
$\mathbf{F}[\hat{\mathbf{q}},\hat{\mathbf{p}}]\!=\!\mathbf{F}\hat{\mathbf{q}}\,\hat{\mathbf{p}}-\mathbf{F}\hat{\mathbf{p}}\,\hat{\mathbf{q}}\!=\!\mathbf{F}\hat{\mathbf{q}}\,\mathbf{F}^{\dagger}\hat{\mathbf{F}}-\mathbf{F}\mathbf{F}^{\dagger}\hat{\mathbf{q}}\mathbf{F}\,\hat{\mathbf{q}}\!=\!-\hat{\mathbf{p}}\,\hat{\mathbf{q}}\mathbf{F}-\hat{\mathbf{q}}\mathbf{F}\,\hat{\mathbf{q}}\mathbf{F}^{\dagger}\mathbf{F}\!=\![\hat{\mathbf{q}},\hat{\mathbf{p}}]\mathbf{F},$
that is
$\mathbf{F}[\hat{\mathbf{q}},\hat{\mathbf{p}}]\!=\![\hat{\mathbf{q}},\hat{\mathbf{p}}]\mathbf{F}.$ (45)
From the relation
[…]
it follows that any set $\mathcal{S}_{\varepsilon}$ is Fourier invariant, that is
$\psi\!\in\!\mathcal{S}_{\varepsilon}\quad\Rightarrow\quad\mathbf{F}\psi\!\in\!\mathcal{S}_{\varepsilon}.$ (47)
Let $\lambda_{0}$, $\lambda_{1}$, …, $\lambda_{d-1}$ be the eigenvalues of $\mathbf{C}\!=\![\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}$, considered in the increasing order of their modulus ($|\lambda_{0}|\leq|\lambda_{1}|\leq...\leq|\lambda_{d-1}|$), and let $\varphi_{0}$, $\varphi_{1}$, … , $\varphi_{d-1}$ be the corresponding eigenfunctions, that is
$\mathbf{C}\varphi_{k}\!=\!\lambda_{k}\ \varphi_{\kappa}.$ (48)
The relation
$\mathbf{C}\varphi_{k}\!=\!\lambda_{k}\,\varphi_{\kappa}\quad\Rightarrow\quad\mathbf{C}\mathbf{F}\varphi_{k}\!=\!\mathbf{F}\mathbf{C}\varphi_{k}\!=\!\lambda_{k}\ \mathbf{F}\,\varphi_{\kappa}$
Table 2: Case d=11d\!=\!11: $\varphi_{1},\,\varphi_{2},...,\varphi_{d}$ are eigenstates of $\mathbf{F}$.
$k$ $\parallel\mathbf{F}\varphi_{k}\!-\!\varphi_{k}\parallel$   $\parallel\mathbf{F}\varphi_{k}\!+\!\varphi_{k}\parallel$   $\parallel\mathbf{F}\varphi_{k}\!-\!\mathrm{i}\varphi_{k}\parallel$   $\parallel\mathbf{F}\varphi_{k}\!+\!\mathrm{i}\varphi_{k}\parallel$
1 $2.0\!\times\!10^{-10}$ 2 1.4142 1.4142
2 1.4142 1.4142 2 $1.2\!\times\!10^{-10}$
3 2 $6.3\!\times\!10^{-11}$ 1.4142 1.4142
4 1.4142 1.4142 $1.3\!\times\!10^{-12}$ 2
5 $2.7\!\times\!10^{-13}$ 2 1.4142 1.4142
6 1.4142 1.4142 2 $2.3\!\times\!10^{-14}$
7 2 $1.3\!\times\!10^{-14}$ 1.4142 1.4142
8 1.4142 1.4142 $3.4\!\times\!10^{-15}$ 2
9 $1.0\!\times\!10^{-15}$ 2 1.4142 1.4142
10 1.4142 1.4142 2 $8.3\!\times\!10^{-16}$
11 2 $8.4\!\times\!10^{-16}$ 1.4142 1.4142
and the numerical data (see Table 2) suggest that the states $\varphi_{0}$, $\varphi_{1}$, … , $\varphi_{d-1}$ are also eigenstates of $\mathbf{F}$, namely Mehta87 ; Barker
$\mathbf{F}\varphi_{n}\!=\!(-\mathrm{i})^{n}\ \varphi_{n}.$ (49)
Particularly, the discrete Fourier transform admits the spectral decomposition
$\mathbf{F}\!=\!\sum_{n=0}^{d-1}(-\mathrm{i})^{n}\ |\varphi_{n}\rangle\langle\varphi_{n}|.$ (50)
For any $\varepsilon>0$ and $d$ high enough, some of the states $\varphi_{0}$, $\varphi_{1}$, … , $\varphi_{d-1}$ belong to $\mathcal{S}_{\varepsilon}$. For example (see Table 1), in the case d=31d\!=\!31, the states $\varphi_{0}$, $\varphi_{1}$, … , $\varphi_{19}$ belong to $\mathcal{S}_{0.001}$.
III.2 A discrete version of the uncertainty relation
A direct consequence of (43) is
$|\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|\approx\mbox{\small$\frac{d}{2\pi}$}\quad\mbox{for}\ \ \psi\!\in\!\mathcal{S}_{\varepsilon}.$ (51)
From the Robertson-Schrödinger uncertainty relation, it follows that the inequality
$\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\frac{|\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi|}{2}\quad\mbox{for}\ \ \psi\!\in\!\mathcal{S}_{\varepsilon},$ (52)
that is
$\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ (53)
is satisfied with approximation in $\mathcal{S}_{\varepsilon}$. See the particular cases presented in the Tables V-XIII and Figure 3.
III.3 Commutation relation of discrete creation-annihilation operators
The operators
[…]
can be regarded as a discrete version of the creation (also called raising) and annihilation (also called lowering) operators
[…]
used in the continuous variable case. For small $\varepsilon\!>\!0$ and any $\psi\!\in\!\mathcal{S}_{\varepsilon}$, we have $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx\mathrm{i}\frac{d}{2\pi}\psi\ \Rightarrow\ [\mathbf{a}^{\dagger},\mathbf{a}]\psi\approx\psi$, that is
[…]
In the continuous-variable case, by using Baker–Campbell–Hausdorff formula, one gets
$\hat{q},\hat{p}]={\rm i}\mbox{\small$\frac{h}{2\pi}$}\quad\Longrightarrow\quad D(q,p)\!=\!\mathrm{e}^{-\frac{\pi\mathrm{i}}{h}qp}\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{h}p\hat{q}}\,\mathrm{e}^{-\frac{2\pi\mathrm{i}}{h}q\hat{p}}\!=\!\mathrm{e}^{\frac{2\pi\mathrm{i}}{h}(p\hat{q}-q\hat{p})}$. (57)
By denoting $\alpha\!=\!\sqrt{\frac{\pi}{h}}(q\!+\!\mathrm{i}p),$ this relation can be written as
$D(\alpha)\!=\!\mathrm{e}^{\alpha\hat{a}^{\dagger}-\bar{\alpha}\hat{a}}$, (58)
It is an open issue if
[…]
IV Some remarkable elements of $\mathcal{S}_{\varepsilon}$ and their properties
IV.1 Discrete-variable Gaussian functions
Figure 1: Gaussian functions $\mathbf{g}_{3}$ and $\mathbf{g}_{1/3}$ in the case d=31d\!=\!31.
For any $\kappa\!\in\!(0,\infty)$, the Gaussian function $g_{\kappa}:\mathbb{R}\rightarrow\mathbb{R}$,
$\begin{array}[]{l}g_{\kappa}(q)\!=\!\mathrm{e}^{-\frac{\kappa\pi}{h}q^{2}},\end{array}$ (60)
satisfies the relation
$F[g_{\kappa}]\!=\!\mbox{\small$\frac{1}{\sqrt{\kappa}}$}\,g_{1/\kappa}.$ (61)
The corresponding discrete-variable Gaussian function $\mathfrak{g}_{\kappa}\!:\!\{-s,\!-s\!+\!\!1,...,s\!-\!\!1,\!s\}\!\rightarrow\!\mathbb{R},$
$\begin{array}[]{l}\mathfrak{g}_{\kappa}(n)=\sum\limits_{\alpha=-\infty}^{\infty}\mathrm{e}^{-\frac{\kappa\pi}{d}(n+\alpha d)^{2}},\end{array}$ (62)
satisfies the relation Vourdas04 ; Cotfas12
$\mathbf{F}[\mathfrak{g}_{\kappa}]\!=\!\mbox{\small$\frac{1}{\sqrt{\kappa}}$}\ \mathfrak{g}_{1/\kappa}$ (63)
similar to (61). It is a particular case of Jacobi $\theta$ function Vourdas04 ; Mehta87 . Because $\mathbf{F}$ is a unitary transform, the corresponding normalized discrete Gaussian functions (see Figure 1)
[…]
satisfy the relation
$\mathbf{F}[\mathbf{g}_{\kappa}]\!=\!\mathbf{g}_{1/\kappa}.$ (65)
For $d$ large enough and small $\delta>0$, there exists $\eta>0$ such that, for $\kappa\!\in\!(\frac{1}{\eta},\eta)$, the most significant coordinates of $\mathbf{g}_{\kappa}$ in the eigenbasis $\{\varphi_{1},\,\varphi_{2},\,...,\,\varphi_{d}\}$ of $\mathbf{C}\!=\![\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}$ are those corresponding to the functions $\varphi_{k}$ with $|\lambda_{k}|<\delta$ (see Table 3). We can consider that
$\hat{\mathbf{q}},\hat{\mathbf{p}}]\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\quad\mbox{for}\ \mathbf{g}_{\kappa}\ \mbox{with}\ \ \kappa\!\in\!(\mbox{\small$\frac{1}{\eta}$},\eta)$ (66)
(see Tables 3, 4 and Figure 2). From (49), we get
$|\langle\varphi_{n}|\mathbf{g}_{\kappa}\rangle|\!=\!|\langle\theta_{n}\varphi_{n}|\mathbf{g}_{\kappa}\rangle|\!=\!|\langle\mathbf{F}\varphi_{n}|\mathbf{g}_{\kappa}\rangle|\!=\!|\langle\varphi_{n}|\mathbf{g}_{1/\kappa}\rangle|,$ (67)
in agreement with the numerical data from Table 3.
Table 3: Case d=11d\!=\!11: for certain $\kappa$, the Gaussian $\mathbf{g}_{\kappa}$ is mainly a linear combination of $\varphi_{1},$ $\varphi_{2},\,\varphi_{3},\,\varphi_{4},\,\varphi_{5}$.
  $\kappa\!=\!1$   $\kappa\!=\!2$   $\kappa\!=\!1/2$   $\kappa\!=\!3$ $\kappa\!=\!1/3$
$|\langle\varphi_{0}|\mathbf{g}_{\kappa}\rangle|$ 0.999968 0.970268 0.970268 0.926811 0.926811
$|\langle\varphi_{1}|\mathbf{g}_{\kappa}\rangle|$ $8.79611\!\times\!10^{-11}$ $8.56453\!\times\!10^{-11}$ $8.50632\!\times\!10^{-11}$ $8.19559\!\times\!10^{-11}$ $8.11208\!\times\!10^{-8}$
$|\langle\varphi_{2}|\mathbf{g}_{\kappa}\rangle|$ $3.14918\!\times\!10^{-11}$ 0.228139 0.228139 0.327299 0.327299
$|\langle\varphi_{3}|\mathbf{g}_{\kappa}\rangle|$ $3.83482\!\times\!10^{-13}$ $3.8548\!\times\!10^{-13}$ $3.58495\!\times\!10^{-13}$ $74436\!\times\!10^{-13} 3.3$$36051\!\times\!10^{-13}$
$|\langle\varphi_{4}|\mathbf{g}_{\kappa}\rangle|$ $0.00798631$ 0.0736351 0.0736351 0.151392 0.151392
$|\langle\varphi_{5}|\mathbf{g}_{\kappa}\rangle|$ $8.61761\!\times\!10^{-15}$ $9.0834\!\times\!10^{-15}$ $8.25462\!\times\!10^{-15}$ $9.15919\!\times\!10^{-15}$ $8.20907\!\times\!10^{-15}$
$|\langle\varphi_{6}|\mathbf{g}_{\kappa}\rangle|$ $2.78121\!\times\!10^{-15}$ 0.0288995 0.0288995 0.0830314 0.0830314
$|\langle\varphi_{7}|\mathbf{g}_{\kappa}\rangle|$ $1.82021\!\times\!10^{-16}$ $1.7344\!\times\!10^{-16}$ $2.47453\!\times\!10^{-16}$ $1.32587\!\times\!10^{-16}$ $3.49103\!\times\!10^{-16}$
$|\langle\varphi_{8}|\mathbf{g}_{\kappa}\rangle|$ $0.000416018$ 0.0146752 0.0146752 0.0549712 0.0549712
$|\langle\varphi_{9}|\mathbf{g}_{\kappa}\rangle|$ $5.0058\!\times\!10^{-17}$ $4.89024\!\times\!10^{-17}$ $6.26338\!\times\!10^{-17}$ $4.68507\!\times\!10^{-17}$ $7.58437\!\times\!10^{-17}$
$|\langle\varphi_{10}|\mathbf{g}_{\kappa}\rangle|$ $1.98563\!\times\!10^{-16}$ 0.00777324 0.00777324 0.0325688 0.0325688
Since $\mathbf{F}$ is a unitary transform, we get
[…]
for any $\psi\!\in\!\mathcal{H}$. Particularly, we have
$\parallel\mathbf{C}\mathbf{g}_{\kappa}\parallel=\parallel\mathbf{C}\mathbf{g}_{1/\kappa}\parallel,$ (69)
in agreement with the numerical data from Table 4.
By using (45) and (65), we get
$\langle\mathbf{g}_{1/\kappa}|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\mathbf{g}_{1/\kappa}\rangle\!=\!\langle\mathbf{F}\mathbf{g}_{\kappa}|\mathbf{F}[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\mathbf{g}_{\kappa}\rangle\!=\!\langle\mathbf{g}_{\kappa}|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\mathbf{g}_{\kappa}\rangle,$ (70)
in agreement with the numerical data from Table 5.
Figure 2: Norm $\parallel([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi})\mathbf{g}_{k}||$ for d=11d\!=\!11 and d=31d\!=\!31.
Table 4: $\hat{\mathbf{q}},\hat{\mathbf{p}}]\mathbf{g}_{\kappa}\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\mathbf{g}_{\kappa}$ is satisfied for certain $\mathbf{g}_{\kappa}$.
$\kappa$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\mathbf{g}_{\kappa}||$
Case $d=11$ Case $d=31$
1 0.0022697 $2.68152\!\times\!10^{-9}$
2 0.283112 0.000670943
1/2 0.283112 0.000670943
3 0.0383181
1/3 0.0383181
Table 5: $|\langle\mathbf{g}_{\kappa}|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\mathbf{g}_{\kappa}\rangle|\approx\frac{d}{2\pi}$ is satisfied for certain $\mathbf{g}_{\kappa}$.
$\kappa$ $|\ |\langle\mathbf{g}_{\kappa}|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\mathbf{g}_{\kappa}\rangle|-\frac{d}{2\pi}\ |$
Case $d=11$ Case $d=31$
1 $1.21005\!\times\!10^{-6}$ $8.88178\!\times\!10^{-16}$
2 0.00350868 $2.47852\!\times\!10^{-9}$
1/2 0.00350868 $2.47852\!\times\!10^{-9}$
3 0.0552824 $7.39351\!\times\!10^{-6}$
1/3 0.0552824 $7.39351\!\times\!10^{-6}$
4 0.204735 0.000383435
1/4 0.204735 0.000383435
5 0.00397916
1/5 0.00397916
6 0.0185731
1/6 0.0185731
7 0.0550503
1/7 0.0550503
8 0.122973
1/8 0.122973
In conclusion, if the dimension $d$ of the Hilbert space describing the quantum system (qudit) is large enough, there exists a family $\mathcal{G}_{\delta}\!=\!\left\{\,\mathbf{g}_{\kappa}\ |\ \kappa\!\in\!(\mbox{\small$\frac{1}{\eta}$},\eta)\ \right\}$ of discrete-variable Gaussian states $\mathbf{g}_{\kappa}$ depending on a continuous parameter $\kappa\!\in\!(\mbox{\small$\frac{1}{\eta}$},\eta)$ such that the relations
$\hat{\mathbf{q}},\hat{\mathbf{p}}]={\rm i}\mbox{\small$\frac{d}{2\pi}$}\quad\mbox{and}\quad\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ (71)
are satisfied with approximation for any $\mathbf{g}_{\kappa}\!\in\!\mathcal{G}_{\delta}$ (See Figure 2 and Figure 3).
Figure 3: Modulus $|\,|\langle\mathbf{g}_{k}|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\mathbf{g}_{k}\rangle|-\frac{d}{2\pi}\ |$ for d=11d\!=\!11 and d=31d\!=\!31.
Any desired precision can be obtained by increasing the dimension $d$ of the Hilbert space and use of an adequate interval $(\mbox{\small$\frac{1}{\eta}$},\eta)$. In addition, the family $\mathcal{G}_{\delta}$ is Fourier invariant:
$\mathbf{g}_{\kappa}\!\in\!\mathcal{G}_{\delta}\quad\Rightarrow\quad\mathbf{F}[\mathbf{g}_{\kappa}]\!=\!\mathbf{g}_{1/\kappa}\!\in\!\mathcal{G}_{\delta}.$ (72)
IV.2 Discrete version of Fock states
It is known that the Hamiltonian of the harmonic oscillator can be written as
$\hat{H}\!=\!\hat{a}^{\dagger}\hat{a}\!+\!\frac{1}{2},$ (73)
and its eigenfunctions are the Fock states
$\Psi_{n}\!=\!\frac{1}{\sqrt{n!}}(\hat{a}^{\dagger})^{n}\Psi_{0}$ (74)
where $\Psi_{0}(q)\!=\!\frac{1}{\sqrt[4]{\pi}}\mathrm{e}^{-\frac{1}{2}q^{2}}$ is the vacuum state. By following the analogy with the continuous case, we define
[…]
where $\mathbf{g}\!\equiv\!\mathbf{g}_{1}$ is the discrete version of the vacuum state.
By direct computation, we can remark that, for small $\varepsilon\!>\!0$ and $d$ high enough, some of the states $\psi_{C,n}$ belong to $\mathcal{S}_{\varepsilon}$. For example (see Table 6), in the case d=31d\!=\!31, the states $\psi_{C,0}$, $\psi_{C,1}$, … , $\psi_{C,8}$ belong to $\mathcal{S}_{0.001}$.
Table 6: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied with approximation for certain states $\psi_{C,k}$.
$\psi$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
$\psi_{C,0}$ $2.68152\!\times\!10^{-9}$ $8.88178\!\times\!10^{-16}$
$\psi_{C,1}$ $3.67539\!\times\!10^{-8}$ 0
$\psi_{C,2}$ $4.93488\!\times\!10^{-7}$ $5.32907\!\times\!10^{-15}$
$\psi_{C,3}$ $2.24984\!\times\!10^{-6}$ $8.88178\!\times\!10^{-15}$
$\psi_{C,4}$ $9.70535\!\times\!10^{-6}$ $2.67342\!\times\!10^{-13}$
$\psi_{C,5}$ 0.000035908 $3.03935\!\times\!10^{-12}$
$\psi_{C,6}$ 0.000125868 $3.62173\!\times\!10^{-11}$
$\psi_{C,7}$ 0.000281649 $1.07889\!\times\!10^{-10}$
$\psi_{C,8}$ 0.000740305 $5.47327\!\times\!10^{-10}$
$\psi_{C,9}$ 0.00153664 $1.43238\!\times\!10^{-9}$
$\psi_{C,10}$ 0.00357292 $1.30984\!\times\!10^{-10}$
IV.3 Discrete version of the harmonic oscillator
The operator $\hat{\mathbf{H}}\!$$\mathcal{H}\!\rightarrow\!\mathcal{H}$,
$\hat{\mathbf{H}}\!=\!\frac{\hat{\mathbf{p}}^{2}\!+\!\hat{\mathbf{q}}^{2}}{2}$ (76)
can be regarded as a discrete version of the Hamiltonian of the quantum harmonic oscillator
$\hat{H}\!=\!\frac{\hat{p}^{2}\!+\!\hat{q}^{2}}{2}.$ (77)
For any small $\varepsilon\!>\!0$, some of the eigenstates of $\hat{\mathbf{H}}$ belong to $\mathcal{S}_{\varepsilon}$. For example (see Table 7), in the case d=31d\!=\!31, the eigenstates $\psi_{O,0}$, $\psi_{O,1}$, … , $\psi_{O,7}$ belong to $\mathcal{S}_{0.001}$. They form an orthonormal system. The relation $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx\mathrm{i}\frac{d}{2\pi}\psi$ is satisfied if and only if $\mathbf{a},\mathbf{a}^{\dagger}]\psi\approx\psi$. In this case $\frac{\hat{\mathbf{p}}^{2}\!+\!\hat{\mathbf{q}}^{2}}{2}\psi\!\approx\!(\hat{\mathbf{a}}^{\dagger}\hat{\mathbf{a}}\!+\!\frac{1}{2})\psi$, and the eigenvalues of $\hat{\mathbf{a}}^{\dagger}\hat{\mathbf{a}}\!+\!\frac{1}{2}$ tend to be distributed equidistant,
$(\hat{\mathbf{a}}^{\dagger}\hat{\mathbf{a}}\!+\!\frac{1}{2})\psi\!=\!\lambda\psi\quad\Longrightarrow\quad(\hat{\mathbf{a}}^{\dagger}\hat{\mathbf{a}}\!+\!\frac{1}{2})\hat{\mathbf{a}}^{\dagger}\psi\!=\!\hat{\mathbf{a}}^{\dagger}(\hat{\mathbf{a}}^{\dagger}\hat{\mathbf{a}}\!+\!\frac{1}{2})\!+\!\hat{\mathbf{a}}^{\dagger}\psi\!=\!(\lambda\!+\!1)\hat{\mathbf{a}}^{\dagger}\psi.$ (78)
Therefore, in the bottom of the spectrum of $\frac{\hat{\mathbf{p}}^{2}\!+\!\hat{\mathbf{q}}^{2}}{2}$, the eigenvalues have an almost equidistant distribution (see Fig. 4).
Figure 4: Eigenvalues of the Hamiltonian (76) in the case d=31d\!=\!31.
Table 7: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied for certain states $\psi_{O,k}$.
$\psi$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
$\psi_{O,0}$ $2.73402\!\times\!10^{-9}$ $5.32907\!\times\!10^{-15}$
$\psi_{O,1}$ $2.31647\!\times\!10^{-8}$ $1.77636\!\times\!10^{-15}$
$\psi_{O,2}$ $3.54883\!\times\!10^{-7}$ $4.44089\!\times\!10^{-15}$
$\psi_{O,3}$ $4.8082\!\times\!10^{-7}$ $1.15463\!\times\!10^{-14}$
$\psi_{O,4}$ $5.27306\!\times\!10^{-6}$ $4.70735\!\times\!10^{-13}$
$\psi_{O,5}$ $0.000018252$ $4.30322\!\times\!10^{-12}$
$\psi_{O,6}$ $0.000167857$ $1.50534\!\times\!10^{-10}$
$\psi_{O,7}$ $0.000127787$ $7.24515\!\times\!10^{-10}$
$\psi_{O,8}$ $0.0010158$ $1.59893\!\times\!10^{-8}$
$\psi_{O,9}$ $0.00233029$ $6.25305\!\times\!10^{-8}$
$\psi_{O,10}$ $0.0176262$ $1.48606\!\times\!10^{-6}$
IV.4 Discrete coherent states
The discrete coherent states Vourdas04 ; Cotfas11 are
$\begin{array}[]{l}|n,k\rangle\!=\!\frac{1}{\sqrt{d}}\mathbf{D}(n,k)|\mathbf{g}\rangle,\end{array}$ (79)
where $n,k\!\in\!\{-\!s,\!-s\!+\!1,...,s\!-\!1,s\}$ and $\mathbf{D}(n,k)\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H},$
$\begin{array}[]{l}\mathbf{D}(n,k)\psi(m)=\mathrm{e}^{-\frac{\pi\mathrm{i}}{d}nk}\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}km}\,\psi(m\!-\!n),\end{array}$ (80)
are the displacement operators. By using the relations
$\parallel\!\mathbf{C}|n,k\rangle\!\parallel^{2}=\!\!\!\sum\limits_{m=-s}^{s}\left|\frac{1}{d}\right.\!\!\sum\limits_{j=-s}^{s}\sum\limits_{\ell=-s}^{s}\!\!j(m\!-\!\ell)\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}j(m-\ell)}\ \mathrm{e}^{\frac{2\pi\mathrm{i}}{d}k\ell}\,\mathbf{g}(\ell\!-\!n)\left.-\,\mathrm{i}\frac{d}{2\pi}\ \mathrm{e}^{\frac{2\pi\mathrm{i}}{d}km}\ \mathbf{g}(m\!-\!n)\right|^{2},$ (81)
$|\bar{z}|\!=\!|z|$, $\mathbf{g}(-n)\!=\!\mathbf{g}(n)$ and the changes of parameters $n\!\mapsto\!-n$, $k\!\mapsto\!-k$, $m\!\mapsto\!-m$, one can prove that
$\parallel\!\mathbf{C}|n,k\rangle\!\!\parallel=\parallel\!\mathbf{C}|-n,k\rangle\!\!\parallel=\parallel\!\mathbf{C}|n,-k\rangle\!\!\parallel=\parallel\!\mathbf{C}|-n,-k\rangle\!\!\parallel.$ (82)
For any small $\varepsilon\!>\!0$, some of the discrete coherent states $|j,k\rangle$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 8), in the case d=31d\!=\!31, the coherent states $|0,0\rangle$, $|0,\pm 1\rangle$, $|0,\pm 2\rangle$, $|0,\pm 3\rangle$, $|0,\pm 4\rangle$, $|\pm 1,0\rangle$, $|\pm 1,\pm 1\rangle$, $|\pm 1,\pm 2\rangle$, $|\pm 1,\pm 3\rangle$, $|\pm 1,\pm 4\rangle$, $|\pm 2,0\rangle$, $|\pm 2,\pm 1\rangle$, $|\pm 2,\pm 2\rangle$, $|\pm 2,\pm 3\rangle$, $|\pm 2,\pm 4\rangle$, $|\pm 3,0\rangle$, $|\pm 3,\pm 1\rangle$, $|\pm 3,\pm 2\rangle$, $|\pm 3,\pm 3\rangle$, $|\pm 3,\pm 4\rangle$, $|\pm 4,0\rangle$, $|\pm 4,\pm 1\rangle$, $|\pm 4,\pm 2\rangle$, $|\pm 4,\pm 3\rangle$, $|\pm 4,\pm 4\rangle$ belong to $\mathcal{S}_{0.001}$.
Table 8: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied with approximation for certain states $|j,k\rangle$.
$|n,k\rangle$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)|n,k\rangle||$   $|\ |\langle n,k|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|n,k\rangle|-\frac{d}{2\pi}\ |$ $|n,k\rangle$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)|n,k\rangle||$ $|\ |\langle n,k|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|n,k\rangle|-\frac{d}{2\pi}\ |$
$|0,0\rangle$ $2.68152\!\times\!10^{-9}$ $8.88178\!\times\!10^{-16}$ $|3,0\rangle$ $8.88946\!\times\!10^{-6}$ $2.70894\!\times\!10^{-13}$
$|0,1\rangle$ $3.53522\!\times\!10^{-8}$ $1.77636\!\times\!10^{-15}$ $|3,1\rangle$ $8.88748\!\times\!10^{-6}$ $2.70894\!\times\!10^{-13}$
$|0,2\rangle$ $6.06924\!\times\!10^{-7}$ $8.88178\!\times\!10^{-16}$ $|3,2\rangle$ $9.21918\!\times\!10^{-6}$ $2.80664\!\times\!10^{-13}$
$|0,3\rangle$ $8.88946\!\times\!10^{-6}$ $2.73559\!\times\!10^{-13}$ $|3,3\rangle$ $0.000011576$ $5.21361\!\times\!10^{-13}$
$|0,4\rangle$ $0.000105793$ $4.01679\!\times\!10^{-11}$ $|3,4\rangle$ $0.000107044$ $4.00791\!\times\!10^{-11}$
$|0,5\rangle$ $0.00102558$ $3.9099\!\times\!10^{-9}$ $|3,5\rangle$ $0.00104948$ $3.91981\!\times\!10^{-9}$
$|1,0\rangle$ $3.53522\!\times\!10^{-8}$ $8.88178\!\times\!10^{-16}$ $|4,0\rangle$ $0.000105793$ $4.01696\!\times\!10^{-11}$
$|1,1\rangle$ $3.53347\!\times\!10^{-8}$ $2.66454\!\times\!10^{-15}$ $|4,1\rangle$ $0.000106023$ $4.01714\!\times\!10^{-11}$
$|1,2\rangle$ $6.2644\!\times\!10^{-7}$ 0 $|4,2\rangle$ $0.000106471$ $4.01306\!\times\!10^{-11}$
$|1,3\rangle$ $8.88748\!\times\!10^{-6}$ $2.74447\!\times\!10^{-13}$ $|4,3\rangle$ $0.000107044$ $4.00791\!\times\!10^{-11}$
$|1,4\rangle$ $0.000106023$ $4.01732\!\times\!10^{-11}$ $|4,4\rangle$ $0.000186876$ $9.42757\!\times\!10^{-11}$
$|1,5\rangle$ $0.00102781$ $3.90986\!\times\!10^{-9}$ $|4,5\rangle$ $0.00100292$ $3.78383\!\times\!10^{-9}$
$|2,0\rangle$ $6.06924\!\times\!10^{-7}$ $1.77636\!\times\!10^{-15}$ $|5,0\rangle$ $0.00102558$ $3.9099\!\times\!10^{-9}$
$|2,1\rangle$ $6.2644\!\times\!10^{-7}$ $8.88178\!\times\!10^{-16}$ $|5,1\rangle$ $0.00102781$ $3.90986\!\times\!10^{-9}$
$|2,2\rangle$ $6.28767\!\times\!10^{-7}$ $2.66454\!\times\!10^{-15}$ $|5,2\rangle$ $0.00103463$ $3.91002\!\times\!10^{-9}$
$|2,3\rangle$ $9.21918\!\times\!10^{-6}$ $2.77112\!\times\!10^{-13}$ $|5,3\rangle$ $0.00104948$ $3.91981\!\times\!10^{-9}$
$|2,4\rangle$ $0.000106471$ $4.01323\!\times\!10^{-11}$ $|5,4\rangle$ $0.00100292$ $3.78383\!\times\!10^{-9}$
$|2,5\rangle$ $0.00103463$ $3.91002\!\times\!10^{-9}$ $|5,5\rangle$ $0.00170788$ $8.59234\!\times\!10^{-9}$
IV.5 Eigenfunctions of discrete annihilation operator
It is known that, in the continuous-variable case, the standard coherent states $|q,p\rangle$ can be defined as eigenstates of the annihilation operator, $\hat{a}|\alpha\rangle\!=\!\alpha|\alpha\rangle$, that is
$\begin{array}[]{l}\hat{a}\,|q,p\rangle\!=\!\sqrt{\frac{\pi}{h}}(q\!+\!\mathrm{i}p)|q,p\rangle.\end{array}$ (83)
By following the analogy, we consider the eigenstates $\psi_{A,n}$ of the discrete version
[…]
of $\hat{a}$. For any small $\varepsilon\!>\!0$, some of the eigenfunctions $\psi_{A,n}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 9), in the case d=31d\!=\!31, the eigenstates $\psi_{A,0}$, $\psi_{A,1}$, … , $\psi_{A,14}$, where the states have been considered in the increasing order of the number of sign oscillations, belong to $\mathcal{S}_{0.001}$.
h $\psi$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$      $\psi$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$ $\psi_{A,0}$ $2.08125\!\times\!10^{-9}$ 0       $\psi_{A,10}$ $3.2588\!\times\!10^{-6}$ $2.75335\!\times\!10^{-14}$ $\psi_{A,1}$ $2.08125\!\times\!10^{-9}$ $8.88178\!\times\!10^{-16}$     $\psi_{A,11}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,2}$ $2.08126\!\times\!10^{-9}$ $1.77636\!\times\!10^{-15}$     $\psi_{A,12}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,3}$ $1.41314\!\times\!10^{-7}$ $8.88178\!\times\!10^{-16}$     $\psi_{A,13}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,4}$ $1.41314\!\times\!10^{-7}$ $8.88178\!\times\!10^{-16}$     $\psi_{A,14}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,5}$ $1.41314\!\times\!10^{-7}$ $1.77636\!\times\!10^{-15}$     $\psi_{A,15}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,6}$ $1.41314\!\times\!10^{-7}$ 0       $\psi_{A,16}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,7}$ $3.2588\!\times\!10^{-6}$ $2.57572\!\times\!10^{-14}$     $\psi_{A,17}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,8}$ $3.2588\!\times\!10^{-6}$ $2.66454\!\times\!10^{-14}$     $\psi_{A,18}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,9}$ $3.2588\!\times\!10^{-6}$ $2.57572\!\times\!10^{-14}$     $\psi_{A,19}$ 0.0381896 $2.72907\!\times\!10^{-6}$
Table 9: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied with approximation by certain states $\psi_{A,k}$.
IV.6 Discrete coherent state quantization
In the continuous case the standard coherent states
$|q,p\rangle\!=\!D(q,p)|g\rangle,$ (85)
where
$\begin{array}[]{l}g(q)\!=\!\frac{1}{\sqrt[4]{\pi}}\,\mathrm{e}^{-\frac{1}{2}q^{2}}\end{array}$ (86)
is the vacuum state, satisfy the decomposition of the identity
$\begin{array}[]{l}\frac{1}{2\pi}\int\!\!\!\int\limits_{\mathbb{R}^{2}}|q,p\rangle\langle q,p|\,dqdp\!=\!\mathbb{I}_{L^{2}(\mathbb{R})}.\end{array}$ (87)
To any function $f\!:\!\mathbb{R}^{2}\!\rightarrow\mathbb{C}$ defined on the phase space $\mathbb{R}^{2}$, and chosen such that the integral is convergent, one can associate the operator
$\begin{array}[]{l}A_{f}\!=\!\frac{1}{2\pi}\int\!\!\!\int\limits_{\mathbb{R}^{2}}f(q,p)\,|q,p\rangle\langle q,p|\,dqdp.\end{array}$ (88)
The map $f\!\mapsto A_{f}$ is called coherent state quantization Cotfas11 ; Gazeau ; Perelomov . One can prove that Gazeau
$\begin{array}[]{lll}A_{f}\!=\!\hat{q}&\mbox{for}&f(q,p)=q,\\ A_{f}\!=\!\hat{p}&\mbox{for}&f(q,p)=p,\\ A_{f}\!=\!\frac{1}{2}(\hat{p}^{2}\!+\!\hat{q}^{2})\!+\!\frac{1}{2}&\mbox{for}&f(q,p)=\frac{1}{2}({p}^{2}\!+\!{q}^{2}).\end{array}$ (89)
Because the discrete coherent states satisfy
$\sum\limits_{n,k=-s}^{s}|n,k\rangle\langle n,k|\!=\!\mathbb{I},$ (90)
by following the analogy with the continuous case, one can associate the operator $\mathbf{A}_{f}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H}$,
$\begin{array}[]{l}\mathbf{A}_{f}\!=\!\sum\limits_{n,k=-s}^{s}f(n,k)\,|n,k\rangle\langle n,k|\end{array}$ (91)
to each function
$f\!:\!\{-s,-s\!+\!1,...,s\}\!\times\!\{-s,-s\!+\!1,...,s\}\!\rightarrow\mathbb{C}$ (92)
defined on the discrete phase space. By using the discrete coherent state quantization $f\!\mapsto\!\mathbf{A}_{f}$, we obtain for $f(n,k)\!=\!\frac{1}{2}(n^{2}\!+\!k^{2})$ the operator
$\begin{array}[]{l}\mathbf{H}_{q}\!=\!-\frac{1}{2}\!+\!\sum\limits_{n,k=-s}^{s}\frac{n^{2}\!+\!k^{2}}{2}\,|n,k\rangle\langle n,k|\end{array}$ (93)
which can be regarded as a discrete version of the harmonic oscillator Hamiltonian. For any small $\varepsilon\!>\!0$, some of the eigenfunctions $\psi_{Q,n}$ of $\mathbf{H}_{q}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 10), in the case d=31d\!=\!31, the states $\psi_{Q,0}$, $\psi_{Q,1}$, … , $\psi_{Q,7}$, where the states have been considered in the increasing order of eigenvalues, belong to $\mathcal{S}_{0.001}$. One can again remark an almost equidistant distribution in the bottom of the spectrum (see Fig. 5). The Hamiltonian (93) is an other discrete version of (77), and a behavior similar to (76) is again expected to happen.
Table 10: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied with approximation of certain states $\psi_{Q,k}$.
$\psi$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
$\psi_{Q,0}$ $3.95539\!\times\!10^{-9}$ $1.77636\!\times\!10^{-15}$
$\psi_{Q,1}$ $3.45774\!\times\!10^{-8}$ $2.66454\!\times\!10^{-15}$
$\psi_{Q,2}$ $5.48854\!\times\!10^{-7}$ $4.44089\!\times\!10^{-15}$
$\psi_{Q,3}$ $6.77894\!\times\!10^{-7}$ $2.4869\!\times\!10^{-14}$
$\psi_{Q,4}$ $7.94567\!\times\!10^{-9}$ $9.82325\!\times\!10^{-13}$
$\psi_{Q,5}$ $0.0000278873$ $8.32134\!\times\!10^{-12}$
$\psi_{Q,6}$ $0.00027139$ $3.52218\!\times\!10^{-10}$
$\psi_{Q,7}$ $0.000187143$ $1.42712\!\times\!10^{-9}$
$\psi_{Q,8}$ $0.00162965$ $3.83563\!\times\!10^{-8}$
$\psi_{Q,9}$ $0.00376062$ $1.33373\!\times\!10^{-7}$
$\psi_{Q,10}$ $0.0310428$ $4.15222\!\times\!10^{-6}$
Figure 5: Eigenvalues of the Hamiltonian (93) in the case d=31d\!=\!31.
IV.7 Mehta states
It is known that, in the continuous-variable case, the eigenstates of the quantum harmonic oscillator $\hat{H}\!=\!\frac{\hat{p}^{2}\!+\!\hat{q}^{2}}{2}$ can be defined in terms of Hermite polynomials as $\Psi_{n}\!:\!(-\infty,\infty)\!\rightarrow\mathbb{R}$,
$\begin{array}[]{l}\Psi_{n}(q)\!=\!\frac{1}{\sqrt{n!\,2^{n}\sqrt{\pi}}}\,H_{n}(q)\ \mathrm{e}^{-\frac{1}{2}q^{2}}\end{array}$ (94)
and satisfy the relation
$\begin{array}[]{l}\hat{H}\Psi_{n}\!=\!(n\!+\!\frac{1}{2})\Psi_{n}.\end{array}$ (95)
It is known that the Hermite-Gauss functions $\Psi_{n}$ are also eigenfunctions of the Fourier transform, namely
$F\Psi_{n}\!=\!(-\mathrm{i})^{n}\Psi_{n}.$ (96)
In his article Mehta87 , Mehta has defined the discrete version $\psi_{M,n}\!:\!\{-s,-s\!+\!1,...,s\}\!\rightarrow\!\mathbb{R}$,
$\begin{array}[]{l}\psi_{M,n}(m)\!=\!\sum\limits_{\alpha=-\infty}^{\infty}H_{n}(m\!+\!\alpha d)\ \mathrm{e}^{-\frac{1}{2}(m\!+\!\alpha d)^{2}}\end{array}$ (97)
of $\Psi_{n}$, and proved that it is an eigenfunction of the discrete Fourier transform,
$\mathbf{F}\psi_{M,n}\!=\!(-\mathrm{i})^{n}\psi_{M,n}.$ (98)
For any small $\varepsilon\!>\!0$, some states $\psi_{M,n}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 11), in the case d=31d\!=\!31, the functions $\psi_{M,0}$, $\psi_{M,1}$, … , $\psi_{M,9}$, after normalization, belong to $\mathcal{S}_{0.001}$.
Table 11: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied with approximation of certain Mehta states $\psi_{M,k}$.
$\psi$   $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
$\psi_{M,0}$ $2.68152\!\times\!10^{-9}$ $8.88178\!\times\!10^{-16}$
$\psi_{M,1}$ $2.41961\!\times\!10^{-8}$ 0
$\psi_{M,2}$ $2.98432\!\times\!10^{-7}$ $2.66454\!\times\!10^{-15}$
$\psi_{M,3}$ $349482\!\times\!10^{-7}$ $1.1543\!\times\!10^{-14}$
$\psi_{M,4}$ $2.95378\!\times\!10^{-6}$ $1.55431\!\times\!10^{-13}$
$\psi_{M,5}$ $8.69175\!\times\!10^{-6}$ $9.28146\!\times\!10^{-13}$
$\psi_{M,6}$ 0.00005669 $1.78559\!\times\!10^{-11}$
$\psi_{M,7}$ 0.00003822 $5.83524\!\times\!10^{-11}$
$\psi_{M,8}$ 0.0002320 $7.42824\!\times\!10^{-10}$
$\psi_{M,9}$ 0.0004216 $2.07663\!\times\!10^{-9}$
$\psi_{M,10}$ 0.0022077 $2.6385\!\times\!10^{-8}$
IV.8 Harper states
In continuous-variable case, $\hat{p}\!=\!F^{\dagger}\hat{q}F$, and the harmonic oscillator Hamiltonian can be written as
$\hat{H}\!=\!\frac{\hat{p}^{2}\!+\!\hat{q}^{2}}{2}\!=\!\frac{\hat{p}^{2}\!+\!F\hat{p}^{2}F^{\dagger}}{2}.$ (99)
A discrete version of the quantum harmonic oscillator can be obtained by using for the differential operator $\frac{\mathrm{d}^{2}}{\mathrm{d}q^{2}}$ the finite-difference version $D^{2}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H}$,
$D^{2}\psi(n)\!=\!\psi(n\!-\!1)\!-\!2\psi(n)\!+\!\psi(n\!+\!1)$ (100)
and the relation
$\begin{array}[]{l}\mathbf{F}D^{2}\mathbf{F}^{\dagger}\psi(n)\!=\!\left(2\cos\frac{2\pi n}{d}\!-\!1\right)\psi(n).\end{array}$ (101)
The eigenfunctions $\psi_{H,n}\!:\!\{-s,-s\!+\!1,...,s,s\!+\!1\}\!\rightarrow\!\mathbb{R}$ of
$\hat{\mathbf{H}}_{d}\!=\!-\frac{1}{2}(D^{2}\!+\!\mathbf{F}D^{2}\mathbf{F}^{\dagger})$ (102)
correspond to the normalized eigenvectors of the matrix (addition is modulo $d$)
$\begin{array}[]{l}\left(2\cos\frac{2\pi n}{d}\delta_{nk}\!-\!4\!+\!\delta_{k,n+1}\!+\!\delta_{k,n-1}\right)_{\!\!-s\leq n,k\leq s}\end{array}$ (103)
and are called Harper functions Barker .
For any small $\varepsilon\!>\!0$, some states $\psi_{H,n}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 12), in the case d=31d\!=\!31, the eigenstates $\psi_{H,0}$, $\psi_{H,1}$, $\psi_{H,2}$ , $\psi_{H,3}$ belong to $\mathcal{S}_{0.001}$. Again, there is an almost equidistant distribution in the lower part of the spectrum (see Fig. 6). The Hamiltonian (102) is an other discrete version of (77), and a behavior similar to (76) is again expected to happen.
Table 12: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied with approximation of certain Harper states $\psi_{H,k}$.
$\psi$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
$\psi_{H,0}$ $1.75165\!\times\!10^{-6}$ $4.08562\!\times\!10^{-14}$
$\psi_{H,1}$ 0.0000101661 $8.56204\!\times\!10^{-13}$
$\psi_{H,2}$ 0.000218006 $1.87562\!\times\!10^{-10}$
$\psi_{H,3}$ 0.000136764 $6.73229\!\times\!10^{-10}$
$\psi_{H,4}$ 0.00211415 $5.87102\!\times\!10^{-8}$
$\psi_{H,5}$ 0.00445289 $1.51256\!\times\!10^{-7}$
$\psi_{H,6}$ 0.0632087 $0.0000150119$
$\psi_{H,7}$ 0.0187208 $0.0000122191$
Figure 6: Eigenvalues of the Hamiltonian (102) in the case d=31d\!=\!31.
IV.9 Kravchuk states
The Kravchuk polynomials $K_{m}$, defined by
$(1\!-\!X)^{s+k}(1\!+\!X)^{s-k}\!=\!\sum\limits_{m=-s}^{s}K_{m}(k)\,X^{s+m}$ (104)
are
$K_{m}(k)\!=\!\sum\limits_{n=0}^{s+m}(-1)^{n}\ C_{s+k}^{m}\ C_{s-k}^{s+m-n}$ (105)
and satisfy the relation
$\begin{array}[]{l}\frac{1}{2^{2s}}\sum\limits_{k=-s}^{s}C_{2s}^{s+k}\ K_{m}(k)\ K_{n}(k)\!=\!C_{2s}^{s+m}\ \delta_{mn},\end{array}$ (106)
where
$C_{m}^{n}\!=\!\left\{\begin{array}[]{cll}\frac{m!}{n!\ (m\!-\!n)!}&\mbox{for}&n\!\in\!\{0,1,...,m\}\\ 0&\mbox{for}&n\!\not\in\!\{0,1,...,m\}.\end{array}\right$. (107)
The Kravchuk functions Nikiforov ; Vilenkin
$\psi_{K,-s},\ \psi_{K,-s+1},\,...\,\ \psi_{K,s}:\{-s,-s\!+\!1,...,s\!-\!1,s\}\!\rightarrow\!\mathbb{R},$
$\psi_{K,m}(n)\!=\!\frac{1}{2^{s}}\sqrt{\frac{C_{2s}^{s+n}}{C_{2s}^{s+m}}}\ K_{m}(n),$ (108)
form an orthonormal basis in $\mathcal{H}$.
For any small $\varepsilon\!>\!0$, some states $\psi_{K,n}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 13), in the case d=31d\!=\!31, the eigenstates $\psi_{K,-15}$, $\psi_{K,-14}$, $\psi_{K,-13}$, belong to $\mathcal{S}_{0.1}$.
Table 13: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied with approximation of certain Kravchuk states $\psi_{K,k}$.
$\psi$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
$\psi_{K,-15}$ 0.00188 $3.3511\!\times\!10^{-8}$
$\psi_{K,-14}$ 0.010478 $1.06622\!\times\!10^{-6}$
$\psi_{K,-13}$ 0.0435235 $0.0000165$
V Some possible physical interpretations
We have proved, mainly numerically, that in the case of a quantum system (qudit) described by a $d$-dimensional Hilbert space, a position-like operator $\hat{\mathbf{q}}$ and a momentum-like operator $\hat{\mathbf{q}}$ satisfy $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx\mathrm{i}\frac{d}{2\pi}\psi$ for certain states $\psi$ if the dimension $d$ is high enough. A fundamental question concerns the existence of any physical interpretation for the operators $\hat{\mathbf{q}}$ and $\hat{\mathbf{p}}$ defined mathematically by following the analogy with the continuous-variable case. It is known that, in the case of a quantum system with $SU(2)$ symmetry, the spectrum of $J_{z}$ has the form $\{-s,-s\!+\!1,..,s\!-\!1,s\}$. So, for example, we can choose $\hat{\mathbf{q}}\!=\!J_{z}$ and $\hat{\mathbf{p}}$ to be the Fourier conjugate operator. More general, any observable with equidistant eigenvalues can play the role of $\hat{\mathbf{q}}$. For eigenvectors forming a basis of the Hilbert space there exists a privileged order, namely the order for which the corresponding eigenvalues form an increasing sequence.
The appearance of quantum states which satisfy with approximation the discrete version of the position-momentum commutation relation among the eigenstates with small energy of certain Hamiltonians suggests a possibility to experimentally obtain such states. They occur naturally when we measure a discrete version of the harmonic quantum oscillator Hamiltonian.
VI Concluding remarks
In the case of a continuous-variable quantum system, the canonical position-momentum commutation relation is exactly satisfied in a subspace of the Hilbert space, and plays a fundamental role. The discrete version of this important relation is also satisfied in a subspace, but not exactly, and this happens only if the dimension of the Hilbert space is high enough. We have investigated the behavior with respect to this discrete version of some important states used in the mathematical description of qudits, including discrete Gaussian states, discrete coherent states and discrete versions of the Hermite-Gauss states. Thus, we have obtained a new property (to our knowledge) of these important quantum states. The discrete version of the position-momentum is satisfied with approximation of certain of them if the dimension of the Hilbert space is high enough. The numerical data presented in this article (Figs. 1-6 and Tables I-XIII) have been obtained using the computer programs presented at the end of the preprint Cotfas-arXiv .
At this moment, there is no satisfactory mathematical explanation for the behavior of the analyzed quantum states concerning the discrete version of the position-momentum commutation relation. Why there exist states for which the relation is satisfied with approximation ? Why they are among the the low energy states ? Why their existence depends on the dimension of the Hilbert space ? There seems to exist a connection between the approximate satisfaction of the position-momentum commutation relation and the tendency of corresponding eigenvalues to be distributed equidistant.
Quantum Information uses a continuous-variable description, but it is mainly limited to a very particular class of states (including the Gaussian states) and a very particular class of unitary transforms (including the Gaussian unitaries). The discrete-variable version of the commutation relation of position -momentum operators is acceptable satisfied by certain particular states. We think that, some useful models involving only such particular states, can be built in the discrete-variable case by following the analogy with the continuous-variable case.
Qudit-based quantum computing Wang2020 is an alternative to the conventional qubit-based quantum computing. The use of quantum systems described by Hilbert spaces of dimension d>2d\!>\!2 provides a larger space to store and process information, and the ability to do multiple control operations simultaneously. These features may allow a reduction of the circuit complexity, a simplification of the experimental setup, and offer the possibility to use more efficient algorithms.